# xGen-MM-Vid (BLIP-3-Video): You Only Need 32 Tokens to Represent a Video Even in VLMs

### v1

@@ 3.2 Public benchmarks
Table 2: Comparison against reported numbers of other models on multiple choice question-answering (MCQ) benchmark.
3.3 Ablations
We conducted an ablation comparing different temporal encoders. These include: (1) the base single frame model (i.e., BLIP-3 trained with videos), (2) mean pooling similar to Video-ChatGPT, and (3) transformer temporal encoder similar to Mirasol3B. We also tried the (4) vanilla Token Turing Machines, which is not the grouped version we use as our temporal encoder.
Table 3 shows the result, comparing the question-answering accuracies of different types of temporal encoders when abstracting a video into 128 tokens. We are able to observe that they all do a reasonable job, while some temporal encoders are more effective.
In addition, we compared different pooling approaches similar to the ones tried in prior works, when they are required to select a much smaller number of tokens (e.g., 32) from a large set of visual tokens. We compare our spatio-temporal attentional pooling as well as the sequential model against its alternatives, including (1) fixed-window (non-learnable) space-time pooling and (2) learnable ‘per-frame’ pooling. In particular, (2) is similar to the approach taken in LLaMA-VID (Li et al., 2024c), which independently selected a fixed number of tokens (e.g., 2) per frame. Table 5 shows the results. Table 5 explicitly compares the impact of having smaller visual tokens. We are able to observe that 32 visual tokens or more gives a reasonable video QA accuracy.
@@ 3.3 Ablations
Ours (Space-time) 77.71
@@
Table 4: Ablations comparing different pooling strategies for 32 tokens.
@@
Speed:
Reducing the number of visual tokens increases the computational efficiency of the models, as the total computation is quadratic to the number of tokens fed to the LLM. We measure the runtime of our models in the training setting for the fair comparison. Here, we report ‘samples per second per GPU’. Without the temporal encoder (i.e., directly using 1024 visual tokens), the model processed 3.3 samples per second. With 16/32/128 tokens using the temporal encoder, the model was able to process 8.5 / 8.2 / 7.5 samples per second.
@@
3.4 Video captioning evaluation
@@
Table 6: Video caption evaluation results using 8 frames. We employ VideoChatGPT’s LLM evaluation and report Average Accuracy / Average Score in this table. The ‘captioning-only model’ was trained only using Mira video caption data (without QA data), making it specialized for the captioning.
@@
We present qualitative video captioning results for the Mira dataset in Figure 4 (more details in Appendix) and for the MSVD and MSRVTT datasets in Figure 5. BLIP-3-Video generates high-quality, detailed captions.
4 Related Works
4.1 Image-Text LLMs
Among recent advances in image-text multimodal models (Li et al., 2023a; Alayrac et al., 2022; Liu et al., 2023; Dai et al., 2023; Xue et al., 2024; Laurençon et al., 2024; Deitke et al., 2024), one common strategy enable image understanding in LLM is to start with a pre-trained image encoder (e.g., ViT (Radford et al., 2021; Zhai et al., 2023)) and a pre-trained language-only LLM (Abdin et al., 2024; Bai et al., 2023; Dubey et al., 2024). The two components are connected via a vision-language connector, which is trained to project vision embeddings output from the vision encoder into “vision tokens" that can be ingested by the LLM. The vision tokens are of the same shape as language embeddings, so the image-text LLM can be trained in the same way as regular language models using the next token prediction loss. There are many design choices for the VL connector, for example, BLIP-2 (Li et al., 2023a) chooses to use a Q-Former to extract vision tokens from the vision embeddings, Flamingo (Alayrac et al., 2022) uses “perceiver resampler" as the connector plus cross-attention layers throughout the language model, while a simpler choice is to use MLP layers to transform the vision embeddings. Image-text LLMs are usually trained with a multi-stage training strategy, including pre-training, instruction tuning, and sometimes, post-training (e.g., DPO (Rafailov et al., 2024)). In addition to simple structured image-text data such as image captioning and single-image VQA, recent works also explore free-from image-text data for model training such as interleaved image-text understanding (Laurençon et al., 2023; Awadalla et al., 2024) and multi-image VQA (Jiang et al., 2024; Li et al., 2024a).
4.2 Video LLMs
Video LLMs extend the architecture of image-based LLMs to handle video input. Zhang et al. (2023) integrates pre-trained encoders and frozen LLMs to process multimodal input through a Video Q-Former and Audio Q-Former, generating video and audio embeddings compatible with LLM without retraining encoders. Maaz et al. (2024) adapts the CLIP visual encoder for video by incorporating temporal features and fine-tunes the model using video-instruction pairs collected by tools like BLIP-2 (Li et al., 2023a) and GRiT (Wu et al., 2022). Li et al. (2024c) generates frame-level embeddings using a visual encoder but condenses visual information into two tokens per frame. However, it does not account for temporal recency across frames. Similarly, models like Video-LLaVA (Lin et al., 2023) and LLaVa-OneVision (Li et al., 2024a) treat videos as long multi-image contexts but lack token efficiency optimization, making them computationally costly. SlowFast-LLaVA (Xu et al., 2024b) adopts a two-stream architecture—Slow and Fast pathways—to capture both spatial and temporal video semantics without extra fine-tuning. Finally, LLaVa-hound-DPO (Zhang et al., 2024b) uses Direct Preference Optimization (DPO) (Rafailov et al., 2024) with GPT-4V to annotate preference data, enhancing video question-answering performance by detecting inconsistencies or hallucinations in model responses.
4.3 Token Pruning
Token pruning is a widely used technique to reduce redundant and overlapping information in Vision Transformers (ViTs) and large language models (LLMs). Bolya et al. (2022) merges similar tokens within ViTs, combining redundant content while retaining task-relevant information across tasks like image, video, and audio processing. Similarly, Ren et al. (2023) employs the Temporal Aggregation Module to combine redundant consecutive video frames and the Spatial Aggregation Module to merge similar patches within each frame, reducing the number of processed tokens by up to 75%. Shen et al. (2024) focus on temporal redundancy and progressively merges tokens across neighboring clips, which reduces the number of tokens by preserving important video-level features. All these methods focus on visual token merging in ViTs, where token processing is challenging in video-based LLMs. In addition, Chen et al. (2024) improves attention efficiency in deeper layers by dynamically pruning or merging redundant image tokens based on attention scores without extra training. Shang et al. (2024) introduces adaptive token reduction through the Adaptive Important Token Selection and Token Supplement, which can be integrated into VLM models without fine-tuning. In LLMs, KV cache pruning is popular for efficient model serving, as seen in (Fu et al., 2024), which uses attention maps to progressively prune tokens and reduce the time-to-first-token (TTFT). Wan et al. (2024) extends KV cache pruning to VLMs, employing different token merging strategies to cut computational costs and support longer multimodal contexts.
@@ 5 Conclusion
Appendix
Figures 6 and 7 compares the video captioning results of BLIP-3-Video against Tarsier and LLaVA-OneVision. The models take the exact same video input with the same number of frames, for a fair comparison. We observe that BLIP-3-Video generate captioning much more similar to the ground truth. It also is able to provide more holistic understanding of the video (rather than focusing on unnecessary details such as texts in the background) while suffering from less hallucinations.

### v2

@@ 3.1 Public benchmarks
[…]
Method (Size) #tokens VC-IT Acc.
PLLaVA (7B) 576+ Y 46.6
VideoLLaMA2 (7B) 1152 Y 54.6
[…]
VideoChat2-Mistral (7B) 96 Y 60.4
Kangaroo (8B) $\sim$10000 Y 61.1
[…]
VideoChatGPT (7B) 264+ N 32.7
VideoLLaMA (7B) 32 N 34.1
VideoChat (7B) 32 N 35.5
LLaMA-VID (7B) 32 N 41.4
Video-LLaVA (7B) 2048 N 43.5
mPLUG-Owl3 (8B) n/a N 54.5
LLaVA-OneVision (7B) 3136 N 56.7
BLIP-3-Video (4B) 32 N 54.9
In addition, we evaluated BLIP-3-Video’s ability to solve multiple choice questions (MCQ). Table 3 shows the results on NExT-QA, and Table 3 shows the results on MVBench. Due to the nature of its questions requiring understanding of multiple frames, many prior models use quite a bit of tokens. For instance, GPT-4 uses a minimum of 255 tokens per frame. It is interesting that BLIP-3-Video achieves comparable accuracy while representing the entire video with only 32 (or 128) tokens.
@@
3.2 Ablations
Temporal encoder: We conducted an ablation comparing different temporal encoders within our model. These include: (1) the base single frame model (i.e., BLIP-3 trained with videos), (2) mean pooling similar to Video-ChatGPT, and (3) transformer temporal encoder similar to Mirasol3B. We also tried (4) the approach of directly extending the spatial Perceiver-Resampler to do spatio-temporal encoding, and (5) vanilla Token Turing Machines, which is not our grouped sequential model.
Table 4 shows the result, comparing the question-answering accuracies of different types of temporal encoders when abstracting a video into 128 tokens. We are able to observe that our sequential model temporal encoder performs the best overall. In particular, directly extending Perceiver-Resampler performed poorly compared to the others.
Table 5: Ablations comparing different pooling strategies (32).
@@ 3.2 Ablations
Ours (Space-time) 77.07
@@
Vs. pooling: We compared against different pooling approaches similar to the ones tried in prior works, when they are required to select a much smaller number of tokens (e.g., 32) from a large set of visual tokens. We compare our spatio-temporal attentional pooling as well as the sequential model against its alternatives, including (1) fixed-window (non-learnable) space-time pooling and (2) learnable ‘per-frame’ pooling. In particular, (2) is similar to the approach taken in LLaMA-VID [27], which independently selected a fixed number of tokens (e.g., 2) per frame. Table 6 shows the results.
Number of visual tokens: Table 6 explicitly compares the impact of having smaller visual tokens. We are able to observe that 32 visual tokens or more gives a reasonable video QA accuracy.
Table 7: Ablations on sequential models (128).
Temporal encoder MSVD-QA ActNet-QA NExT-QA
Original TTM 76.42 / 4.15 54.45 / 3.48 75.42
TTM + time-stamp 76.43 / 4.16 56.15 / 3.53 75.96
TTM + grouped 76.99 / 4.17 55.92 / 3.54 76.46
Ours 77.86 / 4.20 56.66 / 3.56 77.07
Table 8: Ablations on the number of frames
# frames # tokens NExT-QA ActNet-QA
8 frames 32 tokens 76.4 55.7 / 3.5
8 frames 128 tokens 77.1 56.7 / 3.6
16 frames 32 tokens 76.7 55.9 / 3.5
16 frames 128 tokens 77.6 57.3 / 3.6
Grouped token sequential model: Our grouped token sequential model temporal encoder implementation extends the conventional Token Turing Machines. Table 8 shows the ablations confirming the benefits of the new grouped token sequential model formulation. As we observe, the original TTM performs poorly. Adding the time-stamped positional encoding and memory grouping in the sequential model enables much better results, confirming their importance in the real-world models.
More frames: In this experiment we validate whether BLIP-3-Video is able to scale when trained to take more frames. Table 8 shows the results. Even while maintaining the number of tokens, we are able to observe that providing more frames in the input allows BLIP-3-Video to scale to better performance. We believe this is due to the fact that increasing the number of frames has an effect of increasing the size of the “pool” of tokens the temporal encoder can select from. We believe this trend will continue until it saturates.
Speed: Reducing the number of visual tokens increases the computational efficiency of the models, as the total computation is quadratic to the number of tokens fed to the LLM. We measure the runtime of our models in the training setting for the fair comparison. Here, we report ‘samples per second per GPU’. Without the temporal encoder (i.e., directly using 1024 visual tokens), the model processed 3.3 samples per second. With 16/32/128 tokens using the temporal encoder, the model was able to process 8.5 / 8.2 / 7.5 samples per second.
@@
Table 9: Video caption evaluation results using 8 frames. We employ VideoChatGPT’s LLM evaluation and report Average Accuracy / Average Score. The ‘captioning-only model’ was trained only using Mira video caption data (without QA data), making it specialized for the captioning task.
@@
3.3 Video captioning evaluation
@@
We present qualitative video captioning results for the Mira dataset in Figure 4 (and more in Appendix). BLIP-3-Video generates high-quality, detailed captions.
3.4 Longer video datasets
We conducted experiments to confirm BLIP-3-Video’s capability to help long video understanding. We combined BLIP-3-Video’s captioning with a hierarchical text-based video understanding mechanism (i.e., LVNet [32]), and evaluated it on public benchmarks including EgoSchema and VideoMME. We use LVNet [32] as our backbone, relying on its key frame selection by extracting 8 frames centered at those key frames. BLIP-3-Video was applied to such 8-frame segments for each key frame, serving as the video clip captioning model, and the generated captions were given to the high-level LLM in the LVNet framework.
Tables 11 and 11 compares the results. As we observe, compared to the backbone framework used (LVNet), BLIP-3-Video meaningfully improves its accuracy. It captures video information in the short intervals with LVNet, thereby enabling better question answering with long videos. Note that despite BLIP-3-Video was not trained with long videos, it shows promising results.
Table 10: VideoMME long split results.
Method Accuracy (%)
LongVILA 38.8
VideoChat-T 41.9
LLaVA-OneVision 43.8
LLaVA-NexT-Video 44.3
VideoAgent 46.4
Frame-Voyager 51.2
VideoTree 53.1
LVNet 52.4
LVNet + BLIP-3-Video 55.6
LLaVA-Video 61.5
InternVL2.5 62.6
Qwen2-VL 62.2
Table 11: EgoSchema subset results. LVNet* accuracy is obtained by replicating the LVNet results by running its code with GPT-4o as LLM.
Method Accuracy (%)
LLoVi 57.6
VideoAgent 60.2
VideoTree 66.2
LVNet* 66.2
LVNet* + BLIP-3-Video 67.4
LifelongMemory 68.0
Tarsier 68.6
4 Related Work
Image-Text LLMs. Recent image-text multimodal models [24, 2, 29, 8, 51, 21, 9] typically use a pre-trained image encoder (e.g., ViT [34, 54]) and a language-only LLM [1, 5, 10], connected via a vision-language connector that maps visual embeddings into LLM-compatible tokens. These tokens match the shape of language embeddings, allowing joint training via next-token prediction. Connector designs vary: BLIP-2 [24] uses a Q-Former, Flamingo [2] uses a perceiver resampler and cross-attention, and others simply adopt MLPs [29]. Training is often multi-stage—pre-training, instruction tuning, and post-training (e.g., DPO [35])—on both structured (e.g., captioning, VQA) and unstructured data (e.g., interleaved image-text [20, 4], multi-image VQA [15, 22]).
Video LLMs. Video LLMs extend image-based LLMs to handle video input. [56] uses frozen encoders and LLMs with Video/Audio Q-Formers for modality alignment. [31] extends CLIP with temporal features and fine-tunes on video-instruction pairs collected via tools like BLIP-2 [24] and GRiT [46]. [27] compresses frame-level features into two tokens per frame but lacks temporal recency modeling. Video-LLaVA [28] and LLaVa-OneVision [22] treat videos as multi-image sequences, sacrificing compute efficiency. SlowFast-LLaVA [50] introduces dual slow-fast pathways without extra fine-tuning. LLaVa-hound-DPO [57] explores DPO [35] with GPT-4V-generated preference data. Our BLIP-3-Video explores the orthogonal direction of efficiency via reduced tokens.
Token Pruning. Token pruning reduces redundancy in ViTs and LLMs. [6, 36] merges similar tokens within ViTs. [40] focus on temporal redundancy and progressively merges tokens across neighboring clips. [7] prunes or merges tokens in deeper layers based on attention scores. [39] proposes adaptive token reduction via important token selection and supplementation. In LLMs, KV cache pruning improves efficiency [11], and [42] extends this to VLMs with token merging strategies. LongVU [41] spatially compresses frame tokens conditioned on the first frame, orthogonal to temporal reduction. Our BLIP-3-Video utilizes a novel temporal encoder to represent videos in as few as 16 or 32 tokens.
@@ 5 Conclusion
Appendix A Appendix
A.1 More qualitative examples
Figure 5 shows BLIP-3-Video’s captioning results on the MSVD and MSRVTT datasets. We are able to observe that it generates reasonable captions despite utilizing a small number of visual tokens, e.g., 32.
Figures 6 and 7 compares the video captioning results of BLIP-3-Video against Tarsier and LLaVA-OneVision, in detail. The models take the exact same video input with the same number of frames, for a fair comparison. We observe that BLIP-3-Video generate captioning much more similar to the ground truth. It also is able to provide more holistic understanding of the video (rather than focusing on unnecessary details such as texts in the background) while suffering from less hallucinations.
@@ A.1 More qualitative examples
A.2 More experimental results
We tested our model on additional benchmarks: VideoInstruct (which was used in VideoChatGPT) and TempCompass. Table 13 shows the results on the VideoInstruct benchmark. We believe BLIP-3-Video performs very reasonably on all these benchmarks, considering its smaller size and its use of much fewer visual tokens. Similarly, Table 13 shows BLIP-3-Video’s results on ‘Yes/No QA’ and ‘Caption matching’ tasks in TempCompass. They show similar trends to the other experiments. BLIP-3-Video is a compact model giving us a reasonable accuracy.
Table 12: Results on VideoInstruct benchmark.
Model Accuracy
PLLaVA-34B 3.32
SlowFast-LLaVA-34B 3.32
VideoGPT+ 3.28
ST-LLM-7B 3.15
BLIP-3-Video 3.11
VideoChat2_HD_mistral 3.10
LITA-13B 3.04
LLaMA-VID-13B 2.99
VideoChat2 2.98
LLaMA-VID-7B 2.89
Video-ChatGPT 2.38
Table 13: Results on TempCompass benchmark.
Model Yes/No QA Caption matching
GPT-4o 73.66 80.84
Qwen2-VL-7B-Instruct 72.77 77.31
Gemini-1.5-pro 70.32 77.45
LLaVA-OneVision-Qwen-2-7B 69.67 73.79
LLaVA-NeXT-Video-32B-Qwen 69.38 76.51
InternVL2-8B 68.24 77.11
BLIP-3-Video 66.7 66.5
Llama-3-VILA1.5-8B 63.64 68.93
LongVA-7B 62.13 65.67
LLaVA-NeXT-Video-7B-DPO 61.19 63.01
VideoChat2-vicuna-stage3 58.01 53.69
LLaVA-1.5-13B 56.38 64.27
Video-LLaVA-7B 56.38 63.34
Video-LLaMA-2-13B 53.73 54.16
LLaMA-VID-7B-short-video 52.96 56.02
A.3 Reproducibility Statement
We build on top of the open-source BLIP-3 (XGen-MM) model and training code hosted on Huggingface and github. All the experiments were conducted with public datasets. The code and the trained model will be released together with the final version of the paper.
A.4 Limitations
At this point, all the models have been trained with videos up to a couple of minutes. Although we confirmed their capability to handle longer videos by combining it with a framework like LVNet, the model is expected to perform better with better training dataset curation with longer videos. The model sizes (3B) are also relatively small compared to other models. Despite our model showsing very promising model size-accuracy trade-off, training a bigger model with a larger dataset remains as future work.
A.5 Societal Impact
Our proposed BLIP-3-Video significantly improves video-language model efficiency by reducing the number of visual tokens needed question answering. This compact design lowers the computational and energy requirements, enabling more sustainable AI deployments and reducing the environmental footprint of large-scale video understanding. Furthermore, the compact architecture makes high-quality video-language reasoning more accessible to researchers and developers in resource-constrained settings, especially within academia.
As with many powerful AI models, BLIP-3-Video could be misused in ways that raise ethical concerns. The ability to process and understand video content efficiently may be exploited for mass surveillance, privacy-invasive monitoring, or unauthorized profiling. Moreover, if trained on biased or unbalanced data, the temporal encoder may overlook or misrepresent minority-relevant visual cues, potentially reinforcing harmful stereotypes. The model’s compact design could also facilitate easier generation harmful content or misinformation.